OpenStudy (anonymous):

Find the prime factorization of 1170. Please show all of your work.

OpenStudy (anonymous):

factors: 1, 2, 3, 5, 6, 9, 10, 13, 15, 18, 26, 30, 39, 45, 65, 78, 90, 117, 130, 195, 234, 390, 585, 1170 prime factors: 2, 3, 3, 5 ,13

OpenStudy (anonymous):

First step is easy, 10 is a factor, so 1170 = 117 * 10 We know that 10 = 5*2, so 1170 = 117*10 = 117*5*2 117 happens to be divisible by 9 (since the sum of digits is 9). 9 goes in 13 times, so 1170 = 117 * 10 = 117 * 5 * 2 = 9 * 13 * 5 * 2. 9 is 3* 3, so 1170 = 117 * 10 = 117 * 5 * 2 = 9 * 13 * 5 * 2 = 3 * 3 * 13 * 5 * 2

OpenStudy (anonymous):

thank you